a car is at a rate that is 36 mph faster than the rate of a cyclist. the car travels 384 mi in the same amount of time it takes the cyslist to travel 96 mi. find the rate of the car.
Answer by stanbon(75887)

a car is at a rate that is 36 mph faster than the rate of a cyclist.
the car travels 384 mi in the same amount of time it takes the cyslist to travel 96 mi.
find the rate of the car.
----
Cyclist DATA:
dist. = 96 miles ; rate = r mph ; time = d/r = 96/r hrs.
----
Car DATA:
dist = 384 miles ; rate = r+36 mph ; time = d/r = 384/(r+36) hrs.
-------------------------------------------------------
Equation:
time = time
96/r = 384/(r+36)
----
96(r+36) = 384r
384r-96r = 96*36
----
288r = 96*36
r = 12 mph (cyclist's rate)
r+36 = 48 mph (car rate)
=================================
Cheers,
Stan H.
===============
===============================